5 moles = ?????? liters
(1 mole = 22.4 liters)

and how many molecules are in 50 liters?
1
Expert's answer
2010-10-05T19:04:28-0400
5 moles = 22.4*5 = 112 liters;We need to know how many moles are in 50 liters:50/22.4 = 2.23 moles;
2.23 moles * 6.022*1023 mol-1 = 13.44 * 1023 molecules.